OpenStudy (anonymous):

Find a power series for ln(x-1).

zepdrix (zepdrix):

\[\Large\rm \frac{d}{dx}\ln(x-1)=\frac{1}{x-1}\] This implies,\[\Large\rm \ln(x-1)=\int\limits \color{orangered}{\frac{1}{x-1}}dx\]Let's mess with this orange part, turn it into geometric series.

zepdrix (zepdrix):

Factoring out a -1 from the denominator:\[\Large\rm -\int\limits\limits \color{orangered}{\frac{1}{1-x}}dx\]

OpenStudy (anonymous):

Very nice!

zepdrix (zepdrix):

Recall that the convergent geometric series is:\[\Large\rm \sum_{n=0}^{\infty} r^n = \frac{1}{1-r}\]

zepdrix (zepdrix):

So we want to use that in reverse:\[\Large\rm -\int\limits\limits\limits \color{orangered}{\sum_{n=0}^{\infty}x^n}dx\]Make sense so far? :o

zepdrix (zepdrix):

Anyway, from there you can do some fancy business. Bring the integral into the sum,\[\Large\rm -\sum_{n=0}^{\infty} \int\limits x^n ~dx\]Integrating,\[\Large\rm -\sum_{n=0}^{\infty} \frac{1}{n+1}x^{n+1}\]

OpenStudy (anonymous):

I appreciate the help, although i can't really understand it because it's coming through as just about of words and symbols. I can't figure out what you're trying to say! You get an A for effort though :)

OpenStudy (anonymous):

*a bunch of words and symbols

OpenStudy (anonymous):

Apparently it is an issue with my laptop not letting me view it properly. Used my smartphone and can see it just fine. Thank you so much!

OpenStudy (anonymous):

Sometimes the equations do not show up correctly. ln(x-1) is not in the usual table of series you could try to create a Taylor series {check] it out, with derivatives of first and higher orders. #zepdrix showed a clever way to approach the problem recognizing that ln(x-1) is to the integral of dx/(x-1) and that 1/(x-1) is a well-known power series 1 + x + x^2 + x^3..., so the integral gives you x + (1/2) x^2 + (1/3)x^3 + ...

OpenStudy (anonymous):

Yes, I got that I should have the integral of 1/x-1 dx but didn't know where to go from there. So I just say that the power series is x + (1/2)x^2 etc and that's the answer? It can't be that simple can it?????

OpenStudy (anonymous):

I think it is as simple as that: you integrate the 1/(x-1) power series term by trm.